Chapter 8 -  BB exercise No 3 - Practice Question no 1 (1 point)

Mia Corporation manufactured 1,500 units during June. The following variable overhead data

pertain to June:

         Actual variable manufacturing overhead cost                                 $16,800

         Flexible-budget amount for variable manufacturing overhead       $19,000

         Variable manufacturing overhead efficiency variance                    $360 unfavorable

What is the variable overhead flexible-budget variance?

The format of your answer must be as follows:

The variable overhead flexible-budget variance is $10.00 Favorable – (this is not the correct answer)

Chapter 8 -  BB exercise No 3 - Practice Question no 2 (2 points)

Using the data from question 1, answer the following question:

What is the variable overhead spending variance?

         

The format of your answer must be as follows:

The variable overhead spending variance is $10.00 Favorable – (this is not the correct answer)

Chapter 8 -  BB exercise No 3 - Practice Question no 3 (1 point)

Michael Corporation manufactured 26,000 units during March. The fixed-overhead cost-allocation rate is $20.00 per machine-hour. The following fixed overhead data

pertain to March:

                                                                                        Actual              Static Budget

         Production                                                     26,000 units                24,000 units

         Machine-hours                                               6,100 hours                 6,000 hours

         Fixed overhead costs for March                        $128,000                     $120,000

What is the flexible-budget amount?

         

The format of your answer must be as follows:

The flexible-budget amount is $10.00 – (this is not the correct answer)

Chapter 8 -  BB exercise No 3 - Practice Question no 4 (2 points)

Using the data from question 3, answer the following question:

What is the amount of fixed overhead production volume variance?

The format of your answer must be as follows:

The fixed overhead production volume variance is $10.00 Favorable – (this is not the correct answer)
Answer #1

1.variable overhead flexible-budget variance = 16,800 - 19,000 = 2,200 (Favorable)

2. variable overhead spending variance = 2,200 (Favorable) +  $360 (unfavorable) = 2,560 (Favorable)

3.flexible-budget amount = 26,000*(6,000/24,000)*20.00=$130,000

4. fixed overhead production volume variance = 128,000 - 130,000 = 2,000 (Favorable)